PROSPECTS FOR FUNDAMENTAL SYMMETRY TESTS WITH POLYATOMIC MOLECULES

\underline{ROBERT BERGER} and TIMUR ISAEV, Clemens-Schöpf Institute, Technical University of Darmstadt, 64287 Darmstadt, Germany.

Special features of polyatomic molecules make them attractive candidates for search for violation of fundamental symmetries and variation of fundamental constants [1, 2]. We discuss the possibility of searching for nuclear spin-dependent space-parity violating (N SD − PV) interaction in closed-shell and open-shell polyatomic molecules. The parameter W_{a} of the effective molecular spin-rotational Hamiltonian characterising the strength of N SD − PV interaction in open-shell linear molecules is discussed and approaches for its calculation outlined. In addition, possibilities for detecting N SD − PV in chiral molecules via NMR and MW spectroscopy are presented.

REFERENCES:

1. C. Stoeffler et al, Phys. Chem. Chem. Phys., 13 (3), 2011; M. Quack, J. Stohner and M. Willeke, Ann. Rev. Phys. Chem., 59, 2008

2. J. Bagdonaite et al, Science, 339 (6115), 2013.